Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring?
-b b2 - 4ac 2a
Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation:
2x2 + 7x + 3 = 0

Answer by stanbon(75887)

Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring?
-b ;b2 - 4ac ;2a
---
Ans: b^2-4ac
===========================
Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation:
2x^2 + 7x + 3 = 0
-----
b^2-4ac = 49-4*1*3 = 37
===============
Cheers,
Stan H.
===============